20230210 UT


Sunset: 18:23 MST

Evening 12-deg twilight: 19:13 MST

Morning 18-deg twilight: 05:20 MST Morning 12-deg twilight: 05:48 MST

Sunrise: 06:38 MST

Summary

In the afternoon, a critical issue arose with SX M1 (IT 4679). Steve wrote: "While investigating recent instances of IT#2170 (streaming actuator force warnings re: dist board 5), a new serial communication issue arose. The mirror was stuck in panic with various streaming read errors. Spare cards and VxWorks crate were shuttled to the mountain (Thanks DanC, BrianW, SamN). It was determined there was a communication problem with actuator 436 (DanC, JamesR, JohnH, PatH, TimC). Pat and Tim built a null modem adapter to bypass 436 and restore serial communication to the chain. This allowed us to raise the mirror and resume operations on SX."

There was significant cloud cover throughout the night, but DX-only observations were made when it was possible to open. Dave was able to collect data to check the AGw2 focus. There is scatter in the focus curve, but it suggests that the current setting is adequate, and the test allowed him to check and fine tune the procedure for a more thorough measurement in the future.

After Dave's work and some ESM testing, the telescope was closed around 5UT for repair of the SX M1. It was reopened shortly after 10 UT and, because the clouds were still thick and conditions were not optimal for the LBC DDT observations, LBTI optimize gain testing was done.

Jenny's night log, 20230306UT contains the details.

Tasks to be done:

An updated set of LBC DDT targets cover the whole night, so the LBC observations do not need to be restricted to the end of the night. However, the forecast is for cloud cover all night --- we'll keep LBC at the end of the night.

Google doc spreadsheet with outline of plan

https://docs.google.com/spreadsheets/d/1JQxapZ0x6EmD8vzE1nmGI3I3eIBXEFnl/edit?usp=sharing&ouid=110700569789550059640&rtpof=true&sd=true

LBTI twilight - 4:30 UT

  • SW patch
  • optimize gain tests

LUCI AO/ESM - 4:30 (start reconfig) - 10 UT

  • AGw2 focus - LUCI2
  • PSF reconstruction tests - LUCI1
  • SESM - compare med- and high-mode reconstructors - clear and 0.8<seeing<1.5
  • bino-ESM - some clouds and seeing<2"
  • repoint test - any conditions

LBC - 10 (start reconfig) - twilight

  • DDT observation (1 target; 72 deg from moon)
  • photometric standards if clear
  • FPIA injection: to test parameters for Z22 estimate (any clouds, seeing<1.5")
  • thermal aberrations test as a backup (seeing < 1.5", clouds - may give only patches of useful data so may need to be redone)
  • twilight sky flats

SW patches

TCS patches

On Friday, Yang wrote:

I have installed new GCS, OSS, and AOS patches for the TCS system. The GCS and OSS patches are contributed by Paul, and Xianyu authored the new AOS patch. Here is a summary of their purposes:
  1. The GCS patch tries to fix/improve issues reported in IT7752.
  2. The OSS patch addresses IT8827.
  3. The AOS patch adds the dome status to the AOS RTDB.
Note all these new patches are NOT yet activated. The OSA or others will need to manually activate them. Since this Sunday is ECD night, it might be a good time to switch to the new patches. The AOS patch might be needed for some upcoming AO tests. I will leave the decision when to switch and whether to revert after the test to the AO people.

To activate the patches:
  1. ssh tcs@tcs1.mountain.lbto.org
  2. cd /lbt/tcs/current
  3. ./tcs-build latest GCS OSS AOS # or any combinations if you only want to activate some of them
During step 3), it will notify you that it is updating the respective subsystems to ###.20230303 binaries and will ask whether you want to restart the subsystem. Please answer 'y' to the restart question in step 3).

To verify you have the new patches activated:
  1. For GCS, simply launch the GCS GUI, you will see it says "GCS (left/right) 5.7 - GUI 2.1" on the GUI window title bar.
  2. For OSS, on the OSSGUI, if you click Help->About, you should see the OSS version number is 6.4.
  3. For AOS, on the AOSGUI, click the Help->About menu, it should show "18.18" as the AOS version.

In case of problems, you can revert them back (and let the software group know about your issues):
  1. ssh tcs@tcs1.mountain.lbto.org
  2. cd /lbt/tcs/current
  3. ./tcs-build revert GCS OSS AOS # or any combination of them
AO patches

1 active the AOS (can be executed by OSA):

ssh tcs@tcs1.mountain.lbto.org

• cd /lbt/tcs/current

• ./tcs-build latest GCS OSS AOS

1) active the RR mode for AdSec

• ssh -Y AOeng@sxadsec.mountain.lbto.org; ssh -Y

AOeng@dxadsec.mountain.lbto.org

• power off the adsec

• adsc_stop

• use_soul test.20230303.rr

• use_soul; output shows use_soul output shows: Currently

running: test.20230303.rr (soul)

• adsc_start

Note: to test the disable RR mode with the dome status (I can be

available for the test, please give me a call if needed):

• enable_rr_mode.sh

• check the Retroreflector is shown on the AOSGUI

• check the Retroreflector is not shown on the AOSGUI after the

PresetAO command


-- OlgaKuhn - 03 Mar 2023
Topic revision: r6 - 06 Mar 2023, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback